Required: a. Firm D has net income of $32,000, sales of $890,000, and average total assets of $500,000. Calculate the firms margin, turnover, and ROI. (Do not round intermediate calculations. Round your answers to 1 decimal place.) Margin Turnover ROI b. Firm E has net income of $62,000, sales of S1 ,000,000, and ROI of 13%. Calculate the firms turnover and average total assets. (Do not round intermediate calculations. Round Turnover answer to 1 decimal place.) Turnover Average total assets c. Firm F has ROI of 12.6%, average total assets of $1,666,600, and turnover of 1.7. Calculate the firms sales, margin, and net income. (Do not round intermediate calculations. Round Margin answer to 1 decimal place.) Net income Sales Margin
Answer #1

a) calculate following

Margin (32000/890000) 3.6%
Turnover (890000/500000) 1.8 Times
ROI (3.6*1.8) 6.5%

b) Calculate following

Average total assets = 62000/.13 = $476923

Turnover = 1000000/476923 = 2.1 Times

c) Net income = 1666600*12.6% = $209992

Sales = 1666600*1.7 = $2833220

Margin = 209992/2833220 = 7.4%